OpenStudy (anonymous):

Georgio and Maria have decided to carpet 3 rectangular rooms that measure 22 ft by 12 ft, 18 ft by 13 ft, and 14 ft by 10 ft, plus a hallway that is 4 feet by 10 feet. What is the total area to be carpeted, in square feet? How many square yards are needed? Carpet is sold by the square meter. How many square meters of carpet will be needed?

OpenStudy (anonymous):

areas of first room is 22*12 = 264 sq ft same method for the others to convert sq ft to sq yds divide by 9

OpenStudy (anonymous):

Do I just add all the area's together and divide to get the answer in sq yrds? how do I convert to sq m?

OpenStudy (anonymous):

yes to first question to convert to square metres - i think its multiply by 0.836 i'll check up on that

OpenStudy (anonymous):

i got 75.33 sq ft, divided by 9 to get the yr^2 and got 8, and then to get m^2 I got 6.99

OpenStudy (anonymous):

you got 75.33 sq ft as the total? cant be right as the first room area was 264 sq ft

OpenStudy (anonymous):

OH, sorry- so it's 678 sq ft?

OpenStudy (anonymous):

when i divided it to get sq yrds i got 75.33

OpenStudy (anonymous):

first room = 264 2nd = 18*13 = 234 3rd = 140 hall = 40 total = 678 yup divide this by 9 then multiply the result by 0.8361 to get sq m

OpenStudy (anonymous):

yes 75.33 sq yds 75.33 * 0.8361 = 62.99 sq m

OpenStudy (anonymous):

or a nice round 63 sq m

OpenStudy (anonymous):

great! I got confused, because I went on a converting site